OpenStudy (anonymous):

how to find the equation of the line that is perpendicular to y=-3/5x+8 that passes through (0,8)

OpenStudy (anonymous):

Use the formula: y-y1 = m(x-x1), given the m=-3/5

OpenStudy (anonymous):

Plug in your point for y and x

OpenStudy (triciaal):

for perpendicular lines the product of the slope is -1 the slope for the given line is -3/5 so the slope of the perpendicular is -1/(-3/5) = 5/3 this perpendicular line with slope 5/3 passes through the point (0,8) where x = 0 and y = 8 now you can use the formula above to find the eqn of the line